kinzakenny:

A building maintenance company tracked the number of months that fluorescent tubes lasted in the different offices in a building. Calculate the mean, median, and mode for this set of data. 8, 29, 22, 15, 10, 22, 12, 4, 22 Question 7 options: mean 16, median 15, mode, 22 mean 20, median 15, mode, 22 mean 23, median 15, mode, 25 mean 16, median 19, mode, 22

Tranquility:

To calculate the mean, add up all the numbers and divide by how many numbers there are. (There's 9 numbers.) An example: 1, 4, 7, 8 Mean = (1+4+7+8)/4 = 20/4 = 5 To find the median, arrange all the numbers in oder from least to greatest and find the middle number. Ex. 1, 2, 3, 4, 5 3 is the median To find the mode, you need to see which number occurs the most. Ex. 1, 3, 6, 6, 6, 8, 9 5 is the mode because it appears the most times Once you figure it out, let us know so we can check your work. Also welcome to QuestionCove!